# Thermodynamic Signatures of Gaussian Entanglement Beyond Entropy

### v1

@@ Abstract
Continuous-variable quantum thermodynamics in the Gaussian regime provides a promising framework for investigating the energetic role of quantum correlations, particularly in optical systems. In this work, we introduce an entropy-free criterion for entanglement detection in bipartite Gaussian states, rooted in a distinct thermodynamic quantity: ergotropy—the maximum extractable work via unitary operations. By defining the relative ergotropic gap, which quantifies the disparity between global and local ergotropy, we derive two independent analytical bounds that distinguish entangled from separable states. We show that for a broad class of quantum states, the bounds coincide, making the criterion both necessary and sufficient. We further extend our analysis to certain non-Gaussian states and observe analogous energy-based signatures of quantum correlations. These findings establish a direct operational link between entanglement and energy storage, offering an experimentally accessible approach to entanglement detection in continuous-variable optical platforms.
@@
Introduction.– Quantum thermodynamics examines how intrinsically quantum resources shape energy and information processing in microscopic regimes [1, 2, 3, 4, 5, 6, 7, 8, 9]. With experimental advances enabling precise control over single quantum systems, addressing the energetics of nanoscale devices has become both timely and essential [10]. A key concept in this context is ergotropy— the maximum amount of energy that can be extracted from a quantum system through unitary evolution [11]. The importance of this quantity lies in its sole dependency on the Hamiltonian or energy, which essentially captures a distinct, entropy-free notion in the context of quantum resources [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Continuous-variable (CV) systems, particularly quantum optical platforms, offer a compelling setting to explore these thermodynamic aspects. CV states are central to diverse quantum technologies, including quantum communication, sensing, and computation [28, 29, 30]. Among these, Gaussian states are of particular importance due to their experimental feasibility and the efficient implementation of Gaussian operations [31]. These features make them an attractive testbed for probing the energetic role of quantum correlations. Entanglement in Gaussian systems is typically assessed using entropic measures [32] or criteria such as the positive partial transpose (PPT) condition [33]. However, such approaches are grounded in information-theoretic or purely mathematical constructs and do not directly connect to physically measurable energy quantities.
@@
This motivates the development of energy-based entanglement witnesses that operate independently of abstract informational and mathematical metrics. Recent work suggests that ergotropy offers a thermodynamically meaningful framework for revealing quantum correlations [16, 17, 18, 19, 21, 22, 23, 24, 25, 26, 27, 20]. However, these ideas have primarily been explored in discrete-variable (DV) systems. Extending them to CV systems is both conceptually relevant and practically valuable, particularly in the context of emerging quantum technologies, such as quantum batteries, where Gaussian optical systems could provide a scalable and controllable platform [34].
@@
In this work, we introduce the relative ergotropic gap (REG), a quantity that captures the difference in extractable energy between a global bipartite Gaussian state and its locally accessible marginals under Gaussian unitary evolution. We derive analytical bounds that distinguish entangled from separable states, and demonstrate that for a broad class of Gaussian states, REG serves as a necessary and sufficient criterion for entanglement. Operationally, this reveals that entangled states can store and release energy in ways fundamentally beyond the reach of separable states. We further extend our analysis to selected non-Gaussian regimes [35], namely photon-subtracted states, demonstrating that large REG implies higher entanglement (as quantified by the Shchukin-Vogel criterion [36]). These results point toward a broader applicability of energy-based signatures for entanglement in CV systems, with direct relevance to the development of energy-efficient quantum technologies.
@@
Energy Extraction.– In thermodynamics, the amount of work extractable from a system is controlled by its internal energy and entropy. For adiabatic processes, work extraction depends solely on internal energy changes. In the quantum regime, the analogous process is unitary evolution, an isolated dynamics that preserves von Neumann entropy. Ergotropy is the maximum amount of work reversibly extractable [37, 38, 11, 39]. For a state $\rho$ with Hamiltonian $H$, it is defined as $\mathcal{E}(\rho):=E(\rho)-\min_{U}E(U\rho U^{\dagger}),$ where the minimization is over all unitaries $U$, and $E(X):=\operatorname{Tr}(XH)$ denotes the system’s energy in state $X$. This expression quantifies the energy difference between $\rho$ and its passive counterpart $\rho^{p}$, the state with the same spectrum but minimal energy, beyond which no work can be extracted through unitaries. Ergotropy thus quantifies how far a state is from being passive, in purely energetic terms and without reference to entropy.
@@
Energy Extraction in CV Systems.– In systems with infinite-dimensional Hilbert spaces, the implementation of a generic unitary is practically challenging. However, restricting to Gaussian unitaries leads to the concept of Gaussian ergotropy [40, 41], defined as
@@
where $U_{G}$ denotes a Gaussian unitary and $\rho^{p}_{G}$ is the corresponding Gaussian passive state that minimizes the system’s energy under such operations. It is evident that for an arbitrary (possibly non-Gaussian) state $\rho$, Gaussian ergotropy satisfies $\mathcal{E}_{G}(\rho)\leq\mathcal{E}(\rho)$, whereas if $\rho$ is itself a Gaussian state, this inequality becomes an equality because Gaussian unitaries suffice to reach the true passive configuration. In such cases, we omit the subscript G and simply refer to ergotropy.
@@
$\Delta\mathcal{E}_{G}$ $:=\mathcal{E}^{\text{global}}_{G}(\rho)-\mathcal{E}^{\text{local}}_{G}(\rho)=E_{lp}-E_{gp}$ (2)
@@
where $E_{lp}$ and $E_{gp}$ denote the energies of the local and global Gaussian passive states, which are obtained via separable local unitaries $\otimes_{i}U_{i}$ and global Gaussian unitaries $U_{G}$, respectively. Operationally, this quantity captures the energetic advantage of performing correlated (global) operations over local ones. While the notion of ergotropic gap has previously been explored in DV settings as a thermodynamic signature of multipartite quantum correlations [17, 18, 19, 22], its extension to CV systems remains largely unexplored.
Gaussian states constitute the backbone of optical quantum technologies and offer the operational advantage of being completely characterized by their first and second statistical moments. Their simplified mathematical structure enables analytic treatment of thermodynamic quantities such as energy, entropy, and ergotropy, making them a natural playground for quantum thermodynamics in CV systems [29, 40, 41, 42].
@@
In phase space, an $N$-mode bosonic system is described by the quadrature operators $\hat{\vec{r}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})^{T}$, which satisfy the canonical commutation relations $[\hat{r}_{j},\hat{r}_{k}]=i\Omega_{jk}$, where $\Omega=\bigoplus^{N}i\sigma_{Y}$ is the symplectic form [29] and $\sigma_{Y}$ is the Pauli $Y$ matrix.
The energy of a Gaussian state depends on its covariance matrix $\sigma$, with entries $\sigma_{jk}=\langle\{\hat{r}_{j},\hat{r}_{k}\}\rangle-2\langle\hat{r}_{j}\rangle\langle\hat{r}_{k}\rangle$ and the system’s Hamiltonian, which, for a system of quadratically coupled harmonic oscillators with frequencies $\omega_{k}$, can always be brought to the form $\hat{H}=\sum_{k}\omega_{k}(\hat{x}_{k}^{2}+\hat{p}_{k}^{2})$. The average energy of a Gaussian state $\rho$ with covariance matrix $\sigma$ is then
$E(\rho)=\frac{1}{4}\sum_{k}\omega_{k}\left(\text{Tr}[\sigma_{k}]-2\right),$ (3)
@@
where $\sigma_{k}$ is the $2\times 2$ covariance submatrix corresponding to mode $k$. Note that first moments also contribute to the energy of a Gaussian state, but since they can be removed through local unitary operations, they do not affect the ergotropic gap — which quantifies the energy difference between local and global passive states — and thus can be disregarded in our analysis. For a two-mode Gaussian state, one can always bring the covariance matrix to the so-called standard form $\sigma_{sf}$ via local Gaussian unitary operations:
$\sigma_{sf}$ $=\begin{pmatrix}a&0&c_{1}&0\\ 0&a&0&c_{2}\\ c_{1}&0&b&0\\ 0&c_{2}&0&b\end{pmatrix},$ with $\text{with }a\geq b\geq 1\text{ and }c_{1}\geq|c_{2}|.$ (4)
@@
This form is particularly useful for computing both local and global passive energies. The covariance matrix $\sigma_{lp}$ of the local passive state is obtained via local Gaussian operations, while the global passive covariance matrix $\sigma_{gp}$ is achieved via a symplectic transformation that diagonalizes $\sigma_{sf}$ into its Williamson form: $\sigma^{gp}=\text{diag}(\nu_{+},\nu_{+},\nu_{-},\nu_{-})$, where $\nu_{\pm}\geq 1$ are the symplectic eigenvalues of $\sigma_{sf}$, i.e. the moduli of the eigenvalues of $i\Omega\sigma_{sf}$. Assuming $\omega_{A}\leq\omega_{B}$, the ergotropic gap for a two-mode Gaussian state $\rho_{G}$ reads:
$\Delta\mathcal{E}:=E_{lp}-E_{gp}=\frac{1}{2}[\omega_{A}(a-\nu_{+})+\omega_{B}(b-\nu_{-})]$ (5)
@@
where these local and global passive energies are computed directly from the covariance matrices, since they fully determine the passive state’s properties in Gaussian systems.
@@
Main Results: Ergotropic Gap versus Correlations of Gaussian States.– We begin by examining the interplay between quantum correlations and work extraction in the case of pure two-mode Gaussian states. This regime is particularly amenable to analytical treatment and provides a well-understood setting, as the entanglement of pure bipartite states is fully characterized by the entropy of the reduced state.
@@
For pure two-mode Gaussian states $\rho$, the ergotropic gap $\Delta\mathcal{E}(\rho)$ vanishes if and only if $\rho$ is separable. Moreover, $\Delta\mathcal{E}(\rho)$ is a strictly increasing function of the mutual information $I(\rho)$, and hence provides an equivalent quantification of entanglement for these states.
The proof is discussed in Appendix A. Similar to the DV case [18, 19], we find that the ergotropic gap is as a valid entanglement measure for pure two-mode Gaussian states. Since these states are fully characterized by a single parameter, all entanglement measures must be mutually dependent and equivalently capture the quantum correlations. This reflects the situation in pure $2\times d$ dimensional DV systems, where all entanglement measures can be related to the entanglement entropy [43, 44]. The result in Theorem 1 readily extends to pure bipartite Gaussian states of an arbitrary number of modes, by virtue of their mode-wise decomposition into products of two-mode states via local unitaries [45].
The natural question arises regarding whether the ergotropic gap could also serve as a reliable entanglement quantifier for mixed states. Before addressing this, we examine whether $\Delta\mathcal{E}(\rho)$, viewed as a mathematical functional, satisfies basic criteria for a correlation measure [46]. In DV systems, the ergotropic gap has been shown to witness entanglement, but it can be nonzero even for product states [18]. Interestingly, our next result reveals a distinct behavior in the Gaussian setting:
@@
The proof is provided in Appendix D. This establishes that $\Delta\mathcal{E}(\rho)$ acts as a faithful indicator of correlations, being strictly positive if and only if the state is non-product. We now turn to the case of correlated mixed states, exploring whether this quantity can consistently reflect the strength of quantum correlations.
As a preliminary case study, we consider the well-known two-mode squeezed thermal state (TMS), obtained by applying a two-mode squeezing operation with parameter $r$ to two bosonic modes of frequency $\omega$, each initially in thermal equilibrium with a heat bath at temperature $T$. The resulting state’s covariance matrix is itself in standard form:
$\small\sigma_{\text{TMS}}=\left[\begin{matrix}k\cosh{\left(2r\right)}\mathds{1}&k\sinh{\left(2r\right)}\sigma_{z}\\ k\sinh{\left(2r\right)}\sigma_{z}&k\cosh{\left(2r\right)}\mathds{1}\end{matrix}\right],\,\,\text{where}\ \ k=\coth\left(\mbox{$\frac{\omega}{2T}$}\right)\!.$
Its corresponding gap is $\Delta\mathcal{E}_{\text{TMS}}=2k\omega\sinh^{2}r$, which grows monotonically with temperature and diverges as $T\to\infty$. This behavior is particularly striking because thermal noise is generally expected to degrade quantum correlations. Indeed, the quantum mutual information decreases with increasing temperature and remains bounded in the limit of infinite thermal noise: $\lim_{k\rightarrow\infty}I(\sigma_{\text{TMS}})=2\log{\left(\cosh{\left(2r\right)}\right)}$.
This divergence of $\Delta\mathcal{E}$ underscores a fundamental distinction between CV and DV systems. In DV systems, rising temperature leads the state toward a maximally mixed configuration—a global passive state—causing the ergotropic gap to vanish. In contrast, in CV systems, the unbounded energy spectrum allows the thermal population of arbitrarily high energy levels without necessarily rendering the state passive. Consequently, even minimal correlations in extremely energetic states can lead to diverging ergotropic gaps despite diminishing mutual information.
This makes the ergotropic gap $\Delta\mathcal{E}$, in its vanilla form, an unfit candidate to quantify correlations in CV mixed states. To mitigate the divergence problem, we introduce the relative ergotropic gap (REG), defined as
@@
where $E_{gp}$ denotes the energy of the global passive state. This normalized quantity offers a more physically consistent basis for comparison across states. In the following, we investigate whether this quantity can serve as the foundation for a thermodynamically motivated entanglement measure in the CV regime.
To this end we first derive an explicit expression for $\Delta\mathcal{E}_{\text{rel}}$ for a general two-mode Gaussian state. This can be achieved by employing the Bloch-Messiah decomposition [47], which parametrizes general two-mode Gaussian states in terms of elementary Gaussian operations implementable in quantum optical platforms. We consider such operations acting on an initial product state of two thermal modes with temperatures $T_{A}$ and $T_{B}$. The details of the decomposition, along with the derivation of the local and global passive energies, are presented in Appendices C and D. The resulting expression for the REG, as a function of the decomposition parameters, $\Delta\mathcal{E}_{\text{rel}}=f(k,\gamma,\alpha,\theta,z_{1},z_{2})$, is explicitly presented in Appendix D. Here, $k=\frac{k_{A}+k_{B}}{2}$ is the mean thermal fluctuation factor, $\gamma=\frac{k_{A}-k_{B}}{2}$ stands for the fluctuation difference, $\alpha=\frac{\omega_{B}}{\omega_{A}}$ is the ratio between the oscillation frequencies of the two modes, $z_{A}$ and $z_{B}$ are the respective squeezing parameters, and $\theta$ represents the beam splitter angle applied to entangle the modes.
Having introduced the REG, we now reveal a key structural distinction with respect to mutual information:
@@
The relative ergotropic gap $\Delta\mathcal{E}_{\text{rel}}$ is functionally independent of the mutual information.
This independence, demonstrated in Appendix F, highlights the intrinsic significance of the REG and indicates that $\Delta\mathcal{E}_{\text{rel}}$ captures distinct features of quantum correlations not accounted for by mutual information alone. Motivated by this, we explore whether $\Delta\mathcal{E}_{\text{rel}}$ can discriminate between qualitatively distinct types of correlations, such as those in separable versus entangled states.
@@
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$ (7)
(b) If $\rho_{G}$ is entangled, then $\Delta\mathcal{E}_{\text{rel}}(\rho_{G})$ is lower-bounded by
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})>\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)$ (8)
The proof of the above theorem and explicit expressions for the bounds, together with a numerical illustration of the applicability of this result (see Fig.3) are provided in Appendix E. The two bounds are derived analytically employing the Positive Partial Transpose (PPT) criterion [48] as a necessary and sufficient condition for separability in Gaussian two-mode systems. Notably, they depend exclusively on the mean fluctuation parameters of the modes, their fluctuation difference, and the ratio of their respective frequencies. Recall that Lemma 1 establishes $\Delta\mathcal{E}_{\text{rel}}$ as strictly positive for all correlated states — whether classical or quantum — whereas Theorem 2 provides a sharper distinction via two analytic bounds: Eq.7 as a witness for entanglement, and Eq.8 as a witness for separability. These bounds satisfy $\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$, but they do not always coincide. As a result, $\Delta\mathcal{E}_{\text{rel}}$ is a useful qualifier but not a full entanglement quantifier [49], since states in the ‘coexistence region’ [50, 32] characterized by values of $\Delta\mathcal{E}_{\text{rel}}$ between the two bounds cannot be unequivocally classified. However, the bounds do coincide in a physically relevant subclass of states.
@@
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ (9)
@@
The proof of this corollary is provided in Appendix E. As a consequence of these results, the faithfulness of $\Delta\mathcal{E}_{\text{rel}}$ as an entanglement measure is determined by the distance between the bounds $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$: the closer they are, the more accurately our criterion distinguishes between separable and entangled states. In the limiting case where the two bounds coincide, the condition based on the REG becomes equivalent to the PPT criterion. To illustrate this behavior, Fig.1 shows the degree to which $\Delta\mathcal{E}_{\text{rel}}$ violates the separability bound for the parametric family of two-mode squeezed Gaussian states TMS($k,z$), for different values of the fluctuation difference $\gamma$.

### v3

# Ergotropic Characterization of Continuous-Variable Entanglement

@@ Abstract
Continuous-variable quantum thermodynamics in the Gaussian regime provides a promising framework for investigating the energetic role of quantum correlations, particularly in optical systems. In this work, we introduce an entropy-free criterion for entanglement detection in bipartite Gaussian states, rooted in a distinct thermodynamic quantity: ergotropy—the maximum extractable work via unitary operations. By defining the relative ergotropic gap, which quantifies the disparity between global and local ergotropy, we derive two independent analytical bounds that distinguish entangled from separable states. These bounds coincide for a broad class of quantum states, making the criterion both necessary and sufficient in such cases. Unlike entropy-based measures, our ergotropic approach captures fundamentally different aspects of quantum correlations and entanglement, particularly in mixed continuous-variable systems. We also extend our analysis beyond the Gaussian regime to certain non-Gaussian states and observe that Gaussian ergotropy continues to reflect thermodynamic signatures in entangled states, albeit with some limitations. These findings establish a direct operational link between entanglement and energy storage, offering an experimentally accessible approach to entanglement detection in continuous-variable optical platforms.
@@
Introduction.– Quantum thermodynamics examines how intrinsically quantum resources shape energy and information processing in microscopic regimes [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. With experimental advances enabling precise control over single quantum systems, addressing the energetics of nanoscale devices has become both timely and essential [14]. A key concept in this context is ergotropy— the maximum amount of energy that can be extracted from a quantum system through unitary evolution [15]. This makes ergotropy a fundamentally operational and physically measurable quantity, capturing distinct features of quantum states that evade entropy-based analysis [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Continuous-variable (CV) systems, particularly quantum optical platforms, provide a compelling setting for exploring these thermodynamic aspects. CV states are central to diverse quantum technologies, including quantum communication, sensing, and computation [32, 33, 34]. Among these, Gaussian states and operations are particularly important due to their experimental feasibility [35], providing an attractive testbed for probing the energetic role of quantum correlations. Entanglement in Gaussian systems is typically assessed using entropic measures [36] or criteria such as the positive partial transpose (PPT) condition [37, 38]. However, these information-theoretic tools lack direct connections to energy observables and may overlook operational features of quantum correlations.
@@
This motivates the development of energy-based entanglement witnesses that operate independently of abstract mathematical metrics. Recent work suggests that ergotropy offers a thermodynamically meaningful framework for revealing quantum correlations [20, 21, 22, 23, 25, 26, 27, 28, 29, 30, 31, 24, 39]. However, these ideas have primarily been explored in discrete-variable (DV) systems. Extending these concepts to CV systems is both challenging and essential, given their role in practical platforms like quantum batteries and optical processors [40].
@@
In this Letter, we introduce the relative ergotropic gap (REG), which quantifies the disparity between extractable work from a global bipartite Gaussian state and its locally accessible components under Gaussian unitaries. We derive analytical bounds that distinguish separable from entangled states and show that REG is a necessary and sufficient indicator of entanglement in a wide class of Gaussian systems. Crucially, our analysis reveals that for mixed CV states, REG is fundamentally independent of entropic measures, demonstrating an unconventional approach to entanglement characterization. We further extend our energy-based framework to non-Gaussian states [41]. In particular, for photon-subtracted states — where the standard ergotropy is intractable due to their infinite-dimensional support — we use the Shchukin–Vogel criterion [42] to show that REG serves as a reliable entanglement witness in this class. These results highlight the utility of ergotropy as a robust witness of CV quantum correlations and open new directions for energy-driven quantum technologies.
@@
Energy Extraction in CV Systems.– In thermodynamics, the work extractable from a system is controlled by its internal energy and entropy. For adiabatic processes, work extraction depends solely on internal energy changes. In the quantum regime, the analogous process is unitary evolution, which preserves von Neumann entropy. Ergotropy is the maximum amount of work reversibly extractable [43, 44, 15, 45]. For a state $\rho$ with Hamiltonian $H$, it is defined as $\mathcal{E}(\rho):=E(\rho)-\min_{U}E(U\rho U^{\dagger}),$ where the minimization is over all unitaries $U$, and $E(X):=\operatorname{Tr}(XH)$ denotes the system’s energy in state $X$. This expression quantifies the energy difference between $\rho$ and its passive counterpart $\rho^{p}$, the state with the same spectrum but minimal energy, beyond which no work can be extracted through unitaries. Ergotropy thus quantifies how far a state is from being passive, in purely energetic terms.
@@
In systems with infinite-dimensional Hilbert spaces, the implementation of a generic unitary is practically challenging. However, Gaussian states and operations form the backbone of optical quantum technologies, and their simplified mathematical structure makes them a natural playground for quantum thermodynamics in CV systems [33, 46, 47, 48]. Restricting to Gaussian unitaries leads indeed to the concept of Gaussian ergotropy [46, 47], defined as
@@
where $U_{G}$ denotes a Gaussian unitary and $\rho^{p}_{G}$ is the corresponding Gaussian passive state that minimizes the system’s energy under such operations. For an arbitrary state $\rho$, $\mathcal{E}_{G}(\rho)\leq\mathcal{E}(\rho)$, reaching the equality if $\rho$ is itself a Gaussian state, because Gaussian unitaries are optimal; in such cases, we omit the subscript G and simply refer to ergotropy.
@@
$\Delta\mathcal{E}_{G}$ $=\mathcal{E}^{\text{global}}_{G}(\rho)-\mathcal{E}^{\text{local}}_{G}(\rho)=E_{lp}-E_{gp}$ (2)
@@
where $E_{lp}$ and $E_{gp}$ denote the energies of local and global Gaussian passive states. Operationally, this quantity captures the energetic advantage of performing correlated (global) operations over local ones. While the ergotropic gap has been studied in DV settings as a thermodynamic signature of quantum correlations [21, 22, 23, 26], its extension to CV systems remains unexplored.
@@
In phase space, an $N$-mode bosonic system is described by the quadrature operators $\hat{\vec{r}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})^{T}$, which satisfy the canonical commutation relations $\hat{r}_{j},\hat{r}_{k}]=i\Omega_{jk}$, with $\Omega=\bigoplus^{N}i\sigma_{Y}$ the symplectic form [33] and $\sigma_{Y}$ the Pauli $Y$ matrix. The energy of a Gaussian state depends on its covariance matrix $\sigma$, with entries $\sigma_{jk}=\langle\{\hat{r}_{j},\hat{r}_{k}\}\rangle-2\langle\hat{r}_{j}\rangle\langle\hat{r}_{k}\rangle$ and the system’s Hamiltonian, which we assume in the form $\hat{H}=\sum_{k}\omega_{k}(\hat{x}_{k}^{2}+\hat{p}_{k}^{2})$. The mean energy of a Gaussian state $\rho$ with covariance matrix $\sigma$ is then
$E(\rho)=\mbox{$\frac{1}{4}$}{\sum}_{k}\omega_{k}\left(\text{Tr}[\sigma_{k}]-2\right),$ (3)
@@
where $\sigma_{k}$ is the $2\times 2$ covariance submatrix corresponding to mode $k$. Note that first moments also contribute to the energy of a Gaussian state, but since they can be removed through local unitary operations, they do not affect the ergotropic gap and thus can be disregarded in our analysis. For a two-mode Gaussian state, one can always bring the covariance matrix to the standard form $\sigma_{sf}$ via local Gaussian unitary operations:
$\sigma_{sf}$ $=\begin{pmatrix}A&C\\ C&B\\ \end{pmatrix},\ \text{with }A=a\mathbb{I},B=b\mathbb{I},C=\begin{pmatrix}c_{1}&0\\ 0&c_{2}\\ \end{pmatrix}$, (4)
$\ a\geq b\geq 1\text{ and }c_{1}\geq|c_{2}|.$
@@
This form is particularly useful for computing both local and global passive energies. The local passive covariance matrix $\sigma_{lp}$ is obtained via local Gaussian operations, while the global passive covariance matrix $\sigma_{gp}$ is achieved via a symplectic transformation that diagonalizes $\sigma_{sf}$ into its Williamson form: $\sigma_{gp}=\text{diag}(\nu_{+},\nu_{+},\nu_{-},\nu_{-})$, where $\nu_{\pm}\geq 1$ are the symplectic eigenvalues of $\sigma_{sf}$, i.e. the moduli of the eigenvalues of $i\Omega\sigma_{sf}$. Assuming $\omega_{A}\leq\omega_{B}$, the ergotropic gap for a two-mode Gaussian state $\rho_{G}$ reads:
$\Delta\mathcal{E}:=E_{lp}-E_{gp}=\mbox{$\frac{1}{2}$}[\omega_{A}(a-\nu_{+})+\omega_{B}(b-\nu_{-})]$ (5)
@@
where these local and global passive energies are computed directly from the covariance matrices.
@@
Ergotropic Gap versus Correlations of Gaussian States.– We begin by examining the interplay between quantum correlations and work extraction in pure two-mode Gaussian states.
@@
For pure two-mode Gaussian states $\rho$, the ergotropic gap $\Delta\mathcal{E}(\rho)$ vanishes if and only if $\rho$ is separable. Moreover, $\Delta\mathcal{E}(\rho)$ is a strictly increasing function of the quantum mutual information $I(\rho)$, and hence provides an equivalent quantification of entanglement for these states.
The proof is discussed in Appendix B [49]. We find that the ergotropic gap for pure two-mode Gaussian states is fully characterized by a single parameter—just like all standard entanglement measures—keeping them mutually dependent. This observation closely parallels the case of pure $2\times d$ dimensional discrete-variable (DV) systems, where all entanglement measures are functionally related to the entanglement entropy [50, 51]. The result in Theorem 1 readily extends to pure bipartite Gaussian states of an arbitrary number of modes, by virtue of their mode-wise decomposition into products of two-mode states via local unitaries [52]. In contrast, ergotropy-based entanglement measures in higher-dimensional or multipartite DV systems exhibit entropic independence [23, 26], arising from the presence of multiple independent Schmidt coefficients. Interestingly, despite the infinite-dimensional nature of the Gaussian regime, this entropic independence remains absent—highlighting a fundamental structural distinction between Gaussian and general DV systems for pure states.
This leads to a compelling question: can the ergotropic gap also serve as a faithful entanglement measure for mixed states? Moreover, since it arises from closed-system dynamics, does it capture a distinct notion of ‘quantumness’ beyond what is reflected by conventional information-theoretic quantifiers? Before addressing this, we examine whether $\Delta\mathcal{E}(\rho)$ satisfies basic criteria for a correlation measure [53]. In DV systems, the ergotropic gap has been shown to witness entanglement, but it can be nonzero even for product states [22]. Interestingly, our next result reveals a distinct behavior in the Gaussian setting:
@@
The proof is Appendix D.1 [49]. This establishes that $\Delta\mathcal{E}(\rho)$ acts as a faithful indicator of correlations, being strictly positive if and only if the state is non-product. We now turn to the case of correlated mixed states, exploring whether this quantity can consistently reflect the strength of quantum correlations.
As a preliminary case study, we consider the well-known two-mode squeezed thermal state (TMS), obtained by applying a two-mode squeezing operation with parameter $z$ to two bosonic modes of frequency $\omega$, each initially in thermal equilibrium at temperature $T$. From the TMS covariance matrix, reported in Appendix C [49], we find that the ergotropic gap $\Delta\mathcal{E}_{\text{TMS}}=2k\omega\sinh^{2}r$ grows monotonically with temperature and diverges as $T\to\infty$. This is striking because thermal noise is generally expected to degrade quantum correlations. Indeed, the quantum mutual information (QMI) decreases with increasing temperature and remains bounded in the limit of infinite thermal noise: $\lim_{k\rightarrow\infty}I(\sigma_{\text{TMS}})=2\log{\left(\cosh{\left(2r\right)}\right)}$.
This behaviour of $\Delta\mathcal{E}$ underscores a fundamental distinction between CV and DV systems. In DV systems, rising temperature leads the state towards a maximally mixed configuration — a global passive state — causing the ergotropic gap to vanish. In contrast, in CV systems, the unbounded energy spectrum allows the thermal population of arbitrarily high energy levels without necessarily rendering the state passive. Consequently, even minimal correlations in extremely energetic states can lead to large ergotropic gaps despite diminishing QMI. This shows that the ergotropic gap $\Delta\mathcal{E}$, in its ‘vanilla’ form, is not a suitable candidate for quantifying (classical or quantum) correlations in mixed CV states.
To mitigate the divergence problem, we introduce the relative ergotropic gap (REG), defined as
@@
where $E_{gp}$ denotes the energy of the global passive state. First, Lemma 1 guarantees that REG is a faithful signature of (total) correlations. Second, as discussed in Appendix D.2 [49], REG can also be regarded as a faithful witness of more general quantum correlations, like discord. Next, we investigate whether this quantity can capture the finer structure of quantum correlations, namely entanglement in the CV regime.
To this end, we first derive an explicit expression for $\Delta\mathcal{E}_{\text{rel}}$ for a general two-mode Gaussian state. We decompose Gaussian operations acting on an initial product state of two thermal modes with temperatures $T_{A}$ and $T_{B}$ in terms of typical parameters of quantum optical systems [54]. The resulting expression for the REG, as a function of the decomposition parameters, $\Delta\mathcal{E}_{\text{rel}}=f(k,\gamma,\alpha,\theta,z_{1},z_{2})$, is explicitly presented in Appendix D.1 [49]. Here, $k=\frac{k_{A}+k_{B}}{2}$ is the mean thermal fluctuation factor, $\gamma=\frac{k_{A}-k_{B}}{2}$ is the fluctuation difference, $\alpha=\frac{\omega_{B}}{\omega_{A}}$ is the frequency ratio between the two modes, $z_{A}$ and $z_{B}$ are the respective squeezing parameters, and $\theta$ represents the beam splitter angle applied to entangle the modes.
Having introduced the REG, we now reveal a key structural distinction with respect to QMI:
@@
The relative ergotropic gap $\Delta\mathcal{E}_{\textup{rel}}$ is functionally independent of the quantum mutual information.
This independence, demonstrated in Appendix F.1 [49], highlights the intrinsic significance of REG and indicates that $\Delta\mathcal{E}_{\text{rel}}$ captures distinct features of quantum correlations not accounted by QMI alone. To further investigate this discrepancy, we compare entanglement witnesses based on QMI and REG, finding instances where certain entangled states are detected by the latter but not by the former, and vice versa (see Appendix F.2 [49]). Motivated by this, we explore whether $\Delta\mathcal{E}_{\text{rel}}$ can discriminate between qualitatively distinct types of correlations, such as those in separable versus entangled states.
@@
$\Delta\mathcal{E}_{\textup{rel}}(\rho_{G})\leq\mathcal{B}^{\textup{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$ (7)
(b) If $\rho_{G}$ is entangled, then $\Delta\mathcal{E}_{\textup{rel}}(\rho_{G})$ is lower-bounded by
$\Delta\mathcal{E}_{\textup{rel}}(\rho_{G})>\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)$ (8)
The proof of the above theorem and explicit expressions for the bounds, together with a numerical illustration of the applicability of this result, are provided in Appendix E [49]. The two bounds are derived analytically employing the Positive Partial Transpose (PPT) criterion [55] as a necessary and sufficient condition for separability in Gaussian two-mode systems. Notably, they depend exclusively on the mean fluctuation parameters of the modes, their fluctuation difference, and the ratio of their respective frequencies. Recall that Lemma 1 establishes $\Delta\mathcal{E}_{\text{rel}}$ as strictly positive for all correlated states — whether classical or quantum — whereas Theorem 2 provides a sharper distinction via two analytic bounds: Eq. (7) as a witness for entanglement, and Eq. (8) as a witness for separability. These bounds satisfy $\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$, but they do not always coincide. As a result, $\Delta\mathcal{E}_{\text{rel}}$ is a useful qualifier but not a full entanglement quantifier [56], since states in the ‘coexistence region’ [57, 36] characterized by values of $\Delta\mathcal{E}_{\text{rel}}$ between the two bounds cannot be unequivocally classified. However, the bounds do coincide in a physically relevant case.
@@
$\Delta\mathcal{E}_{\textup{rel}}(\rho_{G})\leq\mathcal{B}^{\textup{sep}}_{\mathrm{max}}$ (9)
@@
As a consequence of these results, the faithfulness of $\Delta\mathcal{E}_{\text{rel}}$ as an entanglement measure is determined by the distance between the bounds $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$: the closer they are, the more accurately our criterion distinguishes between separable and entangled states. In the limiting case where the two bounds coincide, the condition based on the REG becomes equivalent to the PPT criterion. To illustrate this behavior, Figure 1 shows the degree to which $\Delta\mathcal{E}_{\text{rel}}$ violates the separability bound for the parametric family of two-mode squeezed Gaussian states TMS($k,z$), for different values of the fluctuation difference $\gamma$.